# Entering Real Social World! Benchmarking the Theory of Mind and Socialization Capabilities of LLMs from a First-person Perspective

### v1

@@ Abstract
In the social world, humans possess the capability to infer and reason about others’ mental states (such as emotions, beliefs, and intentions), known as the Theory of Mind (ToM). Simultaneously, humans’ own mental states evolve in response to social situations, a capability we refer to as socialization. Together, these capabilities form the foundation of human social interaction. In the era of artificial intelligence (AI), especially with the development of large language models (LLMs), we raise an intriguing question: How do LLMs perform in terms of ToM and socialization capabilities? And more broadly, can these AI models truly enter and navigate the real social world? Existing research evaluating LLMs’ ToM and socialization capabilities by positioning LLMs as passive observers from a third-person perspective, rather than as active participants. However, compared to the third-person perspective, observing and understanding the world from an ego-centric first-person perspective is a natural approach for both humans and AI agents. The ToM and socialization capabilities of LLMs from a first-person perspective, a crucial attribute for advancing embodied AI agents, remain unexplored. To answer the aforementioned questions and bridge the research gap, we introduce EgoSocialArena, a novel framework designed to evaluate and investigate the ToM and socialization capabilities of LLMs from a first-person perspective. It encompasses two evaluation environments: static environment and interactive environment, with seven scenarios: Daily Life, Counterfactual, New World, Blackjack, Number Guessing, and Limit Texas Hold’em, totaling 2,195 data entries. With EgoSocialArena, we have conducted a comprehensive evaluation of nine advanced LLMs and observed some key insights regarding the future development of LLMs as well as the capabilities levels of the most advanced LLMs currently available.
@@ 1 Introduction
In the complex social interactions of humans, two fundamental cognitive capabilities play crucial roles: Theory of Mind (ToM) and Socialization. ToM is a fundamental psychological process, defined as the capacity to reason about others’ mental states – beliefs, intents, desires, emotions, knowledge, etc (Premack & Woodruff, 1978; Ma et al., 2023; Chen et al., 2024). Socialization refers to the capability for own mental state evolution in response to social situations. As illustrated in Figure 1(A), When the little boy receives a birthday gift or achieves good grades, he feels very happy. These intertwined capabilities are central to humans’ social life.
With the advent of the era of LLMs, powerful models like GPT-4 (Achiam et al., 2023) and Claude (Anthropic, 2024) have demonstrated remarkable competence in multiple tasks and domains. It raises intriguing questions: How do LLMs perform in terms of ToM and socialization capabilities? And more broadly, can these AI models truly enter and navigate the real social world to achieve efficient human-computer interaction?
To evaluate the ToM and socialization capabilities of LLMs, multiple benchmarks have been proposed, such as SocialIQA (Sap et al., 2022), NormBank (Ziems et al., 2023), BigToM (Gandhi et al., 2023), FanToM (Kim et al., 2023), HI-ToM (Wu et al., 2023), OpenToM (Xu et al., 2024), NegotiationToM (Chan et al., 2024), and ToMBench (Chen et al., 2024). However, as illustrated in Figure 1(B), these benchmarks evaluate LLMs’ ToM and socialization capabilities by positioning LLMs as passive observers from a third-person perspective rather than as active participants. We propose two key points: (1) Observing and understanding the world from an ego-centric first-person perspective is a natural approach for both humans and AI agents. (2) Evaluating LLM’s ToM and socialization capabilities from an ego-centric first-person perspective measures whether LLMs can enter the real social world to achieve efficient human-computer interaction, providing indispensable resources for future research in the field of Embodied Artificial Intelligence (EAI).
Figure 1: (A): ToM (others mental states), Socialization (own mental states). (B): LLM acts as a passive observer to analyze the mental states of characters within a story from a third-person perspective. (C): Evaluation Scenarios in EgosocialArena. (D):Rule-based Agents and RL Agents within Interactive environment.
In this paper, we present EgosocialArena, a novel framework for evaluating and investigating LLMs’ ToM and socialization capabilities from a first-person perspective. We propose a structured taxonomy to guide the development of EgoSocialArena, encompassing three key components: (1) We propose a systematic methodology to transform two existing third-person ToM benchmarks, ToMI and ToMbench, into a first-person perspective. We present a comprehensive argument to justify the validity and rationale behind this transformation. The higher-order ToM questions in a third-person perspective are transformed to evaluate the first-person ToM capabilities of LLMs. In contrast, first-order ToM questions in a third-person perspective are transformed to evaluate LLMs’ mental states in response to social events, i.e., socialization capabilities. Detailed descriptions can be found in 3.1. (2) We investigate the socialization capabilities of LLMs in intriguing and distinctive social situations, including counterfactual, new world, and blackjack game scenarios, as illustrated in Figure 1(C). (3) We investigate the ToM capabilities of LLMs within interactive environments, including belief about the opponent’s behavior patterns in number-guessing scenarios and modeling of opponent’s play styles in Texas Hold’em poker game scenarios. In interactions between models of varying strengths, such as GPT-3.5 and GPT-4, we observe a ”babysitting” issue, where the weaker model can negatively impact the stronger one, diminishing its performance. To mitigate this issue and enable fair evaluation, as illustrated in Figure 1(D), we construct rule-based agents at different cognitive levels and train Reinforcement Learning (RL) agents to provide stable capabilities and behavioral strategies. Overall, EgoSocialArena encompasses two evaluation environments: static environment and interactive environment, with seven scenarios: Daily Life, Counterfactual, New World, Blackjack, Number Guessing, and Limit Texas Hold’em, totaling 2,195 data entries.
We conduct extensive experiments on EgoSocialArena to evaluate 9 foundational models known for their leading performance across multiple tasks and domains. This set includes five API-based models (i.e., o1-preview, GPT-4o, GPT-4-Turbo, GPT-3.5-Turbo, and claude-3-5-sonnet-20240620) and four open-source models (LLaMa-3-8B-Chat, LLaMa-3-70B-Chat, LLaMa-3-8B-Instruct, and LLaMa-3.1-405B-Instruct). We establish a human performance baseline by engaging qualified human annotators. Our experimental results reveal several interesting and critical insights:
  1. 1.
    Although LLMs currently perform behind humans in most scenarios, they have shown significant potential in a few specific cases.
  2. 2.
    The ToM capabilities of LLMs show significant differences between a third-person perspective and a first-person perspective.
  3. 3.
    The powerful capabilities of the o1-preview model are truly surprising.
  4. 4.
    The scaling up of open-source models or performing instruction fine-tuning has not yielded significant results in terms of ToM and socialization capabilities.
  5. 5.
    The performance gap between open-source models and API-based models in socialization scenarios is significant.
2 Related Work
Existing ToM Benchmarks
Previous evaluations for the ToM of LLMs primarily focus on testing models using narrative stories, also referred to as reading comprehension scenarios. Specifically, Le et al. (2019) proposed the ToMi benchmark based on the classic Sally-Anne test. Wu et al. (2023) introduced the HI-ToM benchmark, which focuses on higher-order belief reasoning and sets up scenarios where agents can communicate with each other. Gandhi et al. (2023) proposed BigToM, which presents a framework for designing a ToM benchmark from synthetic templates for evaluating different aspects of LLMs’ ToM capabilities (e.g., desire and belief). Xu et al. (2024) introduced OpenToM, which assigns personalities to agents in the stories and ensures that the storylines are more reasonable and logical. Chen et al. (2024) proposed ToMBench, which systematically evaluates LLMs across all dimensions of ToM capabilities. Unlike the above methods that require LLMs to read stories and answer related questions, some studies evaluate LLMs’ performance by inputting dialogues to them. Kim et al. (2023) proposed FanToM, which tests LLMs on their ability to infer the mental states of characters in everyday conversations. Chan et al. (2024) introduced NegotiationToM, which restricts the dialogue content to negotiation scenarios.
Existing Socialization Benchmarks
Sap et al. (2022) proposed SocialIQA and used it to evaluate LLMs. SocialIQA contains many questions related to social commonsense. Ziems et al. (2023) introduced NormBank, a large repository of social norms knowledge, which can be used to assess social norm-related tasks. Li et al. (2024) reorganized and classified existing datasets related to social intelligence. Xu et al. (2023) studied LLMs’ understanding of the world and explored how different persuasion strategies could modify LLMs’ worldviews. Different from our work, the aforementioned ToM and socialization benchmarks evaluate LLMs by placing them in a passive observer’s third-person perspective.
Necessity of developing LLMs’ ToM and socialization capabilities
Several studies (Gandhi et al., 2023; Kim et al., 2023) have shown that LLMs have poor reasoning performance and robustness on ToM and socialization tasks in a zero-shot setting, even with the current state-of-the-art GPT series (Achiam et al., 2023) model. With LLMs becoming increasingly integrated into our everyday lives, developing LLMs with ToM and socialization could be better at teaching us, learning from us, communicating with us, collaborating with us, and understanding us (Gandhi et al., 2021; 2023; Rabinowitz et al., 2018; Shu et al., 2021).
Ego-centric (First-person Perspective) Research
In the fields of computer vision and robotics, there has already been considerable research on a first-person perspective. For example, Cheng et al. (2023) explored whether vision-language models can ”Think from a First-person Perspective?” Huang et al. (2023) proposes the construction of embodied agents in a 3D world, which involves acquiring and processing first-person perspective images. Huang et al. (2024) built a bridge between third-person and first-person perspectives at the action level, while Dou et al. (2024) proposed a method designed to transform exocentric video-language data for egocentric video representation learning. However, research on first-person perspectives in the field of natural language processing remains unexplored.
@@ 2 Related Work
LLMs for Interactive Scenarios
Some work focuses on designing interaction strategies to enable LLMs to gain more benefits during interactions. For example, Zhang et al. (2024a) proposed Agent-pro, Zhang et al. (2024b) introduced K-level reasoning, and Guo et al. (2023) put forward the Suspicion-Agent. Additionally, Li et al. (2023) explored Multi-LLM collaboration by informing LLMs of task rules through prompts. Park et al. (2023) introduced generative agents that can simulate human behavior. Bianchi et al. (2024) explored the social behavior of LLMs in negotiation scenarios. Fu et al. (2023) show LLMs can improve each other in a negotiation scenario. Fan et al. (2024) examined the capability of LLMs to make rational decisions in game theoretic scenarios.
3 EgoSocialArena
EgoSocialArena provides a systematic approach to convert third-person perspective ToM benchmarks into first-person perspective, constructs Rule-based Agents at different cognitive levels and Reinforcement Learning (RL) Agents, and includes various scenarios to evaluate LLMs’ ToM and socialization capabilities from a first-person perspective.
3.1 Converting Existing Third-person ToM Benchmarks to a First-person Perspective
Foundation and Inspiration
In LLM-based Agent applications, system message serves as a critical component, functioning to pre-set the model’s role and background. As illustrated in Figure 2(A), system message ”You are name and live in a town…” is used. Interestingly, in the domain of LLM self-awareness research (Laine et al., 2024), a similar linguistic construct is employed. As illustrated in Figure 2(B), researchers employ the pronoun ”you” to probe LLMs’ potential self-awareness. Inspired by and building upon studies in these two domains, we systematically modify system message, story, question, and answer options to transform third-person ToM benchmarks into a first-person perspective.
@@ 3.1 Converting Existing Third-person ToM Benchmarks to a First-person Perspective
As illustrated in Figure 2(C), unlike instructing LLMs in system message that ”you are a helpful assistant.”, we inform LLMs in system message that they have personally experienced certain social events, similar to deploy LLM-based Agent. As illustrated in Figure 2(D), we employ the pronoun ”you” to replace specific characters in stories and questions, thereby situating LLMs within particular roles. This approach enables the models to experience social events from a first-person perspective. The framing of questions is akin to that employed in self-awareness research. For modifications to answer options, consider LLMs answer from a first-person perspective, substituting ’I’ for specific character in the options. Higher-order ToM questions from third-person perspective benchmarks, after conversion, are still used to evaluate the ToM capabilities of LLMs from a first-person perspective. In contrast, first-order ToM questions, after conversion, form an assessment of LLMs’ mental states following social events, which we include in the scope of evaluating socialization capabilities.
3.2 Intriguing and Distinctive Social Situations
As illustrated in Figure 3, we design three particularly interesting scenarios—Counterfactual, New World, and Blackjack—all used to evaluate the socialization capabilities of LLMs.
Counterfactual
In real social situations, the rules of Rock-Paper-Scissors are: rock beats scissors, scissors beat paper, and paper beats rock. An LLM can relatively easily establish a belief based on this situation. In contrast, we define a counterfactual situation for the Rock-Paper-Scissors game (scissors beat rock, paper beats scissors, and rock beats paper) to explore whether an LLM can establish a belief that matches this counterfactual situation.
Figure 3: Evaluation examples for the Counterfactual, New World, and Blackjack scenarios in intriguing and distinctive social situations, as well as Number Guessing scenarios in the interactive environment.
New World
We design stories that describe new social world scenarios that are significantly different from the current social world. We aim to investigate whether LLMs can demonstrate cognitive flexibility and achieve substantive shifts in their understanding and reasoning about the social world.
Blackjack
In the single-turn card game scenario, it is necessary to analyze the game situation, i.e., analyze own cards in Self-belief and the opponent’s cards in External-belief. This forms a comprehensive mental estimation of the current game situation. We experiment with the blackjack card game, and its specific rules can be found in A.1.
3.3 Interactive Environment
In the interactive environment, many studies adopt two different LLMs to interact (e.g., GPT-3.5 vs GPT-4). However, we have observed a phenomenon in such experiments: ”Babysitting” weaker model distracts the stronger model. When comparing the ToM capabilities of the Deepseek-v2 model and the GPT-3.5 model from a first-person perspective, it is unreasonable to conduct experiments like GPT-3.5 vs GPT-4 and Deepseek-v2 vs GPT-4. Although both opponents are GPT-4, the babysitting phenomenon during the interaction process significantly interferes with the opponent’s output, leading to unfair comparisons. To ensure a fair and reasonable comparison, the opponent’s capability and behavioral strategies must be controlled to be stable during the interaction process. We propose two methods: constructing rule-based agents at different cognitive levels as opponents and training RL agents as opponents.
3.3.1 Rule-based Agents at Different Cognitive Levels
@@
Level 1: $a_{o}^{t}=C$. In this pattern, we conduct experiments with the rule-based Agent’s actions remaining constant at 50. Level 2: $a_{o}^{t}=f(t)=50-5(t-1)$. In this pattern, we conduct experiments with the rule-based Agent’s action sequence of round 1: 50, round 2: 45, …, round 9: 10, round 10: 5, an arithmetic sequence with the first term 50 and a common difference of 5. Level 3: $a_{o}^{t}=f(a_{m}^{t-1},a_{o}^{t-1})=0.8\times\left(\frac{a_{m}^{t-1}+a_{o}^{t-1}}{2}\right)$. In this pattern, we conduct experiments with the rule-based Agent’s action copying the gold value from the previous round. The rules of number guessing can be found in A.1.
3.3.2 Reinforcement Learning Agents
@@
In the Limit Texas Hold’em scenario, we train two reinforcement learning agents as opponents: Deep Q-network (DQN)-Aggressive (Mnih et al., 2015) and DQN-Conservative (Mnih et al., 2015). By adapting the reward function, RL agents are given different game personalities. For DQN-Aggressive, we encourage the action of raising and calling during the game. In contrast, for DQN-Conservative, we encourage the action of folding during the game. The rules of Limit Texas Hold’em can be found in A.1, and a specific example of the Limit Texas Hold’em scenario can be found in A.2.
3.3.3 Interaction
During the interaction between an LLM and a rule-based agent in Number Guessing scenario, as illustrated in Figure 3, each round will involve asking the LLM $\mathcal{M}$ ToM questions $\mathcal{Q}_{ng}$ regarding the opponent’s behavior:
$\mathcal{A}_{ng}^{t}=\mathcal{M}(\mathcal{Q}_{ng}^{t}\oplus\mathcal{H}^{\textless t}\oplus\mathcal{P}_{ng}),t\in\{1,2,\dots,N\}$ (1)
where $\mathcal{P}_{ng}$ is a prompt for leading to the answer, $t$ represents the current round number, <$\mathcal{H}^{\textless t}$ represents game history information, $\mathcal{A}_{ng}^{t}$ represents LLM’s response.
After multiple rounds of game interaction between the LLM and the RL Agent in Limit Texas Hold’em scenario, asking the LLM $\mathcal{M}$ ToM questions $\mathcal{Q}_{lth}$ regarding the opponent’s play style:
$\mathcal{A}_{lth}^{t}=\mathcal{M}(\mathcal{Q}_{lth}^{t}\oplus\mathcal{H}^{\textless t}\oplus\mathcal{P}_{lth}),t=N$ (2)
where $\mathcal{P}_{lth}$ is a prompt for leading to the answer, $t$ represents the current round number, <$\mathcal{H}^{\textless t}$ represents game history information, $\mathcal{A}_{lth}^{t}$ represents LLM’s response.
4 Data construction
4.1 Data Collection and Validation

### v2

# Entering Real Social World! Benchmarking the Social Intelligence of Large Language Models from a First-person Perspective

@@ Abstract
Social intelligence is built upon three foundational pillars: cognitive intelligence, situational intelligence, and behavioral intelligence. As large language models (LLMs) become increasingly integrated into our social lives, understanding, evaluating, and developing their social intelligence are becoming increasingly important. While multiple existing works have investigated the social intelligence of LLMs, (1) most focus on a specific aspect, and the social intelligence of LLMs has yet to be systematically organized and studied; (2) position LLMs as passive observers from a third-person perspective, such as in Theory of Mind (ToM) tests. Compared to the third-person perspective, ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios. (3) a lack of comprehensive evaluation of behavioral intelligence, with specific emphasis on incorporating critical human-machine interaction scenarios. In light of this, we present EgoSocialArena, a novel framework grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence, aimed to systematically evaluate the social intelligence of LLMs from a first-person perspective. With EgoSocialArena, we conduct a comprehensive evaluation of eight prominent foundation models, even the most advanced LLMs like o1-preview lag behind human performance by 11.0 points. Our code and data are available at https://github.com/gyhou123/EgoSocialArena.
@@ 1 Introduction
Social intelligence, i.e., the ability to understand and reason about the mental states of others (cognitive intelligence), awareness and adaptation to the social context (situational intelligence), and effective interaction with others (behavioral intelligence), is a form of advanced intelligence that naturally develops during human growth Thorndike (1921); Hunt (1928); Premack and Woodruff (1978); Hou et al. (2024); Li et al. (2024). Imagine the future where robots powered by large language models (LLMs) enter our social world, communicating with us empathetically, supporting us in living better, and making great contributions to society. This is a wonderful vision and highlights the importance and significance of understanding, evaluating, and developing the social intelligence of LLMs.
Numerous datasets have been curated to assess the social intelligence of LLMs, such as ToMI Le et al. (2019), BigToM Gandhi et al. (2023), FanToM Fan et al. (2024), HI-ToM Wu et al. (2023), OpenToM Xu et al. (2024), and ToMBench Chen et al. (2024b) for evaluating Theory of Mind (ToM) capabilities of LLMs, focusing on reasoning about the mental states of others; SocialIQa Sap et al. (2022) and NormBank Ziems et al. (2023) for evaluating LLMs’ understanding of social contexts; SOTOPIA Zhou et al. (2023) and LLMArena Chen et al. (2024a) for evaluating LLMs’ behavior and interaction capabilities in social goal-driven and gaming scenarios. However, as illustrated in Figure 1(A), these existing works each focus on a specific aspect of social intelligence, such as ToM tests corresponding to cognitive intelligence, and the social intelligence of LLMs has not yet been systematically organized and studied.
On the other hand, as illustrated in Figure 1(B), these existing works evaluate LLMs’ ToM and social context understanding abilities by positioning LLMs as passive observers from a third-person perspective. We propose two key points: (1) The third-person perspective involves making LLMs engage in "armchair theorizing" that isn’t aligned with real LLM-based Agent use scenarios. This kind of evaluation isn’t accurate enough. (2) Ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios, allowing us to better and more thoroughly understand their performance in human society.
Figure 1: (A): Datasets related to social intelligence over time in the Era of LLMs (a non-exhaustive visualization due to space constraints). (B): LLM acts as a passive observer to analyze mental states of characters within a story from a third-person perspective. (C): Main direction of existing work on the behavioral intelligence of LLMs.
Moreover, as illustrated in Figure 1(C), when evaluating the behavioral and interactive capabilities of LLMs, existing work like LLMArena propose various game environments and have different LLMs interact to see who wins and loses. Compared to having two LLMs play games to determine winners and losers, exploring LLM’s performance in human-machine interaction is more meaningful. Additionally, many works, such as Hypothetical Minds Cross et al. (2024) and SOTOPIA-Pi Wang et al. (2024), focus on proposing various strategies, such as prompt-based methods or behavior cloning, to enhance the performance of LLMs in interactive environments like Melting 2.0 Agapiou et al. (2022) and SOTOPIA. However, there is still a lack of comprehensive evaluation of behavioral intelligence for current mainstream LLMs.
In this paper, we present EgoSocialArena, a novel framework designed to systematically evaluate the social intelligence of LLMs from a first-person perspective. The development of EgoSocialArena is grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence: (1) Cognitive Intelligence: we propose a complete and generalizable workflow to transform existing static third-person ToM benchmarks into a first-person perspective. Additionally, we have newly developed a dynamic cognitive assessment in multi-turn interactive scenarios. (2) Situational Intelligence: Imagine an LLM-based Agent entering our social world - how would it respond emotionally when receiving praise or gifts [^1] This might be related to self-awareness, but the focus could be shifted more towards the application situations.? We have newly developed an assessment for such real-world social situations. Additionally, we have also developed assessments for counterfactual situations and parallel world situations. (3) Behavioral Intelligence: we incorporate existing cooperative and adversarial game environments, as well as social goal-driven interactive dialogue environments, to comprehensively evaluate the behavioral intelligence of LLMs. Overall, as illustrated in Figure 2, EgoSocialArena encompasses the evaluation of cognitive, situational, and behavioral intelligence, with eight scenarios: static cognition, dynamic cognition evolution, real-world social situation, counterfactual situation, parallel world situation, cooperative game, adversarial game, and social goal-driven human-machine interactive dialogue environment, comprising a total of 2245 data entries.
Figure 2: Examples of eight scenarios in EgoSocialArena.
We conduct extensive experiments on EgoSocialArena to evaluate 8 foundational models known for their leading performance across multiple tasks and domains. This set includes five API-based models (i.e., o1-preview, GPT-4o, GPT-4-Turbo, GPT-3.5-Turbo, and claude-3-5-sonnet-20240620) and three open-source models (LLaMa-3-8B-Chat, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-Instruct). We establish a human performance baseline by engaging qualified human annotators with a college degree or higher. Our experimental results reveal several interesting and critical insights: (1) The o1-preview model achieved the highest score of 80.6 among all models, surpassing human performance in dynamic cognition and adversarial game scenarios. Nevertheless, an 11.0 gap in overall accuracy remains when compared to the human baseline, leaving plenty of room for model improvement. Our in-depth analysis reveals that the superiority of o1-preview is mainly attributed to its powerful logical reasoning and mathematical abilities (keenly uncovering deeper patterns behind the data). (2) Comparing the performance of LLaMA3-8B-Chat and LLaMA3-70B-Chat models shows that simply scaling model size does not significantly help improve the social intelligence of LLMs. (3) Compared to the third-person perspective, LLMs show significantly improved ToM reasoning ability when operating from a first-person perspective.
2 EgoSocialArena
2.1 Cognitive Intelligence
In the static cognition scenario, we convert the existing third-person ToMI benchmark to a first-person perspective. In the dynamic cognition evolution scenario, we construct opponents with various behavioral strategies, including rule-based agents at different cognitive levels and Reinforcement Learning (RL) agents, to explore how LLMs can form beliefs about opponents’ behavioral strategies during multi-round interactions.
2.1.1 Static Cognition — Converting Existing Third-person ToM Benchmarks to a First-person Perspective
Foundation and Inspiration
In LLM-based Agent applications, system message serves as a critical component, functioning to pre-set the model’s role and background. As illustrated in Figure 3(A), system message "You are name and live in a town…" is used. Interestingly, in the domain of LLM self-awareness research (Laine et al., 2024), a similar linguistic construct is employed. As illustrated in Figure 3(B), researchers employ the pronoun "you" to probe LLMs’ potential self-awareness. Inspired by and building upon studies in these two domains, we systematically modify system message, story, question, and answer options to transform third-person ToM benchmarks into a first-person perspective.
@@ 2.1.1 Static Cognition — Converting Existing Third-person ToM Benchmarks to a First-person Perspective
As illustrated in Figure 3(C), unlike instructing LLMs in system message that "you are a helpful assistant.", we inform LLMs in system message that they have personally experienced certain social events, similar to deploy LLM-based Agent. As illustrated in Figure 3(D), we employ the pronoun "you" to replace specific characters in stories and questions, thereby situating LLMs within particular roles. This approach enables the models to experience social events from a first-person perspective. The framing of questions is akin to that employed in self-awareness research.
2.1.2 Dynamic Cognition Evolution — Number Guessing (G0.8A)
Scenario: G0.8A
Each player selects a number between 1 to 100. The objective is to select a number that is closest to 80% of the group’s average choice.
Rule-based Agents at Different Cognitive Levels
@@
Level 1: $a_{o}^{t}=C$. In this pattern, we conduct experiments with the rule-based Agent’s actions remaining constant at 50. Level 2: $a_{o}^{t}=f(t)=50-5(t-1)$. In this pattern, we conduct experiments with the rule-based Agent’s action sequence of round 1: 50, round 2: 45, …, round 9: 10, round 10: 5, an arithmetic sequence with the first term 50 and a common difference of 5. Level 3: $a_{o}^{t}=f(a_{m}^{t-1},a_{o}^{t-1})=0.8\times\left(\frac{a_{m}^{t-1}+a_{o}^{t-1}}{2}\right)$. In this pattern, we conduct experiments with the rule-based Agent’s action copying the gold value from the previous round.
2.1.3 Dynamic Cognition Evolution — Limit Texas Hold’em
Scenario: Limit Texas Hold’em
The game commences with each player being dealt two private cards Five community cards are then dealt face-up in a series of stages: a three-card Flop, followed by a single card on the Turn and another single card on the River. The player can choose from four actions: Fold, Check, Call, Raise.
Reinforcement Learning Agents
@@ 2.1.3 Dynamic Cognition Evolution — Limit Texas Hold’em
In the Limit Texas Hold’em scenario, we train two reinforcement learning agents as opponents: Deep Q-network (DQN)-Aggressive (Mnih et al., 2015) and DQN-Conservative (Mnih et al., 2015). By adapting the reward function, RL agents are given different game personalities. For DQN-Aggressive, we encourage the action of raising and calling during the game. In contrast, for DQN-Conservative, we encourage the action of folding during the game. A specific example of the Limit Texas Hold’em scenario can be found in Appendix B.
2.2 Situational Intelligence
2.2.1 Real-World Social Situation
By filtering data from SocialIQa and ToMBench and using the transformation method mentioned in section 2.1.1, we evaluate the mental states of LLMs’ self after experiencing certain social events from a first-person perspective.
2.2.2 Counterfactual Situation
The conventional rules of Rock-Paper-Scissors (RPS) are: rock beats scissors, scissors beat paper, and paper beats rock. An LLM can relatively easily adapt to this situation. In contrast, we define a counterfactual situation for the RPS game (scissors beat rock, paper beats scissors, and rock beats paper) to explore whether an LLM can achieve situational adaptation. In addition to constructing counterfactual situations like RPS games, we also construct counterfactual situations based on physical facts, chemical facts, biological facts, traffic rules, social etiquette knowledge, etc.
2.2.3 Parallel World Situation
We design narratives depicting parallel social world that differ significantly from our current social world. We aim to investigate whether LLMs can demonstrate situational adaptation to these alternative worlds.
2.3 Behavioral Intelligence
2.3.1 Adversarial Game
Blackjack, also known as 21 points, is a card game that involves a dealer and a player. The player must decide whether to hit or stand based on own hand, the dealer’s face-up card, and the dealer’s one hidden card. The objective is to beat the dealer without exceeding 21 points. We evaluate the win rate of LLMs as a player in this scenario.
2.3.2 Cooperative game
Defuse Bomb: Three LLMs emulate specialists in a team to defuse bombs. Bombs are distributed across $n$ rooms, whether the rooms are interconnected can be set manually. Each bomb exhibits unique phase sequences in $m$ colors, requiring the correct order of wire cutters for defusing. Team members start with different colored cutters and must coordinate and synchronize efforts for efficiency. We create 5 different map environments, each containing 5 bombs. Following Li et al. (2023), each successfully defused bomb awards the team 10 points per processed phase. We measure collaboration efficiency by calculating the score a team composed of three LLMs can achieve within 10 rounds.
2.3.3 Social-goal Driven Human-Machine Interactive Dialogue
With an open-ended social interaction environment SOTOPIA Zhou et al. (2023), which assigns a social goal and character to each agent involved. We focus on a comprehensive evaluation of interactions between current mainstream LLMs and humans, aiming to provide a more intuitive comparison of behavioral differences between humans and LLMs in social goal-driven interactive dialogue. We use the goal completion metric to quantitatively express this difference.
3 Data Collection, Validation and Statistics